OpenStudy (nincompoop):

my favorite calc 1 question: The "Folium of Descartes" has the following equation, \( x^3+y^3-6xy = 0 \) a) sketch the tangent line at the point \( (\frac{4}{3},\frac{8}{3}) \) b) Find the slope of the tangent line to the graph at \( (\frac{4}{3},\frac{8}{3}) \) c) Find the equation of the tangent line at the point \( (\frac{4}{3}, \frac{8}{3}) \)

OpenStudy (nincompoop):

besides figuring out what the graph looks like, the rest are easy to solve using: implicit differentiation (b), and point-slope form (c).

OpenStudy (nfcfox):

WAIT WAAAAAAAAAAT

OpenStudy (nincompoop):

http://youtu.be/jG4DYZ5uuE0

OpenStudy (nincompoop):

The "Folium of Descartes" has the following equation, \( x^3+y^3-6xy = 0 \) a) sketch the tangent line at the point \( (\frac{4}{3},\frac{8}{3}) \) b) Find the slope of the tangent line to the graph at \( (\frac{4}{3},\frac{8}{3}) \) c) Find the equation of the tangent line at the point \( (\frac{4}{3}, \frac{8}{3}) \)

OpenStudy (nincompoop):

b) Find the slope of the tangent line to the graph at \( (\frac{4}{3},\frac{8}{3}) \) \( \frac{d}{dx}(x^3+y^3-6xy) = \frac{d}{dx}0 \\ 3x^2+3y^2 y' -6(y+xy') = 0 \\ 3x^2 + 3y^2y' - 6y - 6xy' = 0 \\ 3y^2y'-6xy'=-3x^2+6y \\ y'(3y^2-6x) = -3x^2+6y \\ \sf \large y' = \frac{-3x^2+6y}{3y^2-6x} \\ clean~up \\ \large y' = \frac{-3(x^2-2y)}{-3(-y^2+2x)} \rightarrow y'=\frac{x^2-2y}{-y^2+2x} \) plug-in the coordinate \( \sf \large(\frac{4}{3}, \frac{8}{3}) \) \[ y' =\large \frac{(\frac{4}{3})^2-2(\frac{8}{3})}{-(\frac{8}{3})^2+2(\frac{4}{3})} \rightarrow \large \frac{\frac{16}{9}-\frac{16}{3}}{-\frac{64}{9}+\frac{8}{3}}\\ \large \frac{\frac{-32}{9}}{\frac{-40}{9}} \rightarrow \frac{-32}{-40} \rightarrow \color {blue}{\frac{4}{5}} \] c) Find the equation of the tangent line at the point \( (\frac{4}{3}, \frac{8}{3}) \) \( y-y_1 = m(x-x_1) \\ y \large - \frac{8}{3} = \frac{4}{5}(x-\frac{4}{3}) \\ y = \large \frac{4}{5}x-\frac{16}{15}+\frac{40}{15} \\ \color{blue}{y = \large \frac{4}{5}x+\frac{24}{15}} \)

OpenStudy (anonymous):

So weird...

OpenStudy (xapproachesinfinity):

so how does the graph look like!

OpenStudy (xapproachesinfinity):

to sketch the tangent line i need to sketch the graph of that equation first

OpenStudy (nincompoop):